Integrate dx/(a-x)(b-x)=k2dt in second order rate of a reaction and explain th integration

Grade:

1 Answers

Pankaj
9 years ago

Integrated rate law for a zero order reaction is :

[R]=-kt +[R]0

When we compare the equation with the equation of a straight line i.e. y = mx + c,

we get slope of line = m = -k and Y intercept = [R]0

Slope is equal to –k for the plot [R] vs t

Same equation can also be written as

t = 1/k([Ro]-[R])

When compared with the equation of straight line y = mx + c

The slope would be 1/k for the plot t vs ([Ro]-[R]) and the y intercept would be 0 as c = 0

Hope it’s clear to you.

Thanks & Regards

Pankaj Singh

askIITians Faculty